# Lattice isomorphic Banach lattices of polynomials

### v1

@@ 3 Regular nuclear polynomials
Let $w\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ be given. By [4, Lemma 3.5], there exist a finite dimensional vector sublattice $F_{1}$ of $F$ and a vector $w_{1}\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\otimes F_{1}$ such that $\|w-w_{1}\|_{|\pi|}\leq 2^{-1}$ and $\|w_{1}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{1}}\leq 2\|w_{1}\|_{|\pi|}+1$. Using [4, Lemma 3.5] again, there exist a finite dimensional vector sublattice $F_{2}$ of $F$ and a vector $w_{2}\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\otimes F_{k}$ such that $\|(w-w_{1})-w_{2}\|_{|\pi|}\leq 2^{-2}$ and $\|w_{2}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{2}}\leq 2\|w_{2}\|_{|\pi|}+2^{-1}$. Proceeding inductively, we may construct a sequence $(F_{k})_{k}$ of finite dimensional vectors sublattices of $F$ and a sequence $(w_{k})_{k}\subset\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ such that $w_{k}\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\otimes F_{k}$, $\|(w-w_{1}-\cdots-w_{k-1})-w_{k}\|_{|\pi|}\leq 2^{-k}$ and $\|w_{k}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq 2\|w_{k}\|_{|\pi|}+2^{1-k}$ for every $k>1$. Thus,
$\left|\|w\|_{|\pi|}-\sum_{i=1}^{k}\|w_{i}\|_{|\pi|}\right|\leq\|w-\sum_{i=1}^{k}w_{i}\|_{|\pi|}\leq 2^{-k}$
@@
holds for every $k\in\mathbb{N}$. By letting $k\to\infty$, we get that $\|w\|_{|\pi|}=\sum_{i=1}^{\infty}\|w_{i}\|_{|\pi|}<\infty$ and $w=\displaystyle\sum_{k=1}^{\infty}w_{k}$. We also notice that $\|w_{k+1}\|_{|\pi|}\leq\|w\|_{|\pi|}-\sum_{i=1}^{k}\|w_{i}\|_{|\pi|}\leq 2^{-k}$ holds for every $k\in\mathbb{N}$, and so
$\|w_{k}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{k}}$ $\leq 2\|w_{k}\|_{|\pi|}+2^{1-k}\leq 2\,2^{-(k-1)}+2^{1-k}$
$=2\,2^{-k+1}+2^{1-k}=2^{1-k}(2+1)=\frac{6}{2^{k}}$
@@
$\|w_{1}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{1}}\leq 2\|w_{1}\|_{|\pi|}+1\leq 2\|w\|_{|\pi|}+1\leq 3=\displaystyle\frac{6}{2^{1}}$
@@
If $F$ is a Dedekind complete Banach lattice with the LAP, then the map $\overline{\Phi}_{n}\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;F)$ is injective, and hence an isometry.
@@
Suppose for the sake of contradiction that $\overline{\Phi}_{n}$ is not injective. Thus, there exists $w\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ with $\|w\|_{|\pi|}=1$ such that $\overline{\Phi}_{n}(w)=0$. Let $(F_{k})_{k}$ and $(w_{k})_{k}$ be as in Lemma 3.1. Since, for every $k\in\mathbb{N}$, $w_{k}\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\otimes F_{k}$ and $\|w_{k}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq 2\|w_{k}\|_{|\pi|}+2^{1-k}\leq\displaystyle\frac{6}{2^{k}}$, there are a positive elements $z_{k,1},\dots,z_{k,m_{k}}\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}$ and positive vectors $y_{k,1},\dots,y_{k,m_{k}}\in F_{k}$ such that
$|w_{k}|\leq\displaystyle\sum_{j=1}^{m_{k}}z_{k,j}\otimes y_{k,j}\quad\text{ and}\quad\displaystyle\sum_{j=1}^{m_{k}}\|z_{k,j}\|_{s,|\pi|}\|y_{k,j}\|\leq\frac{6}{2^{k}}.$
For each $k\in\mathbb{N}$ and $j=1,\dots,m_{k}$, we let
$u_{k,j}=k\|y_{k,j}\|z_{k,j}\quad\text{and}\quad v_{k,j}\displaystyle=\left\{\begin{matrix}\frac{y_{k,j}}{k\|y_{k,j}\|},\,\text{if }y_{k,j}\neq 0;\\ 0,\,\text{if }y_{k,j}=0.\end{matrix}\right$.
Thus, $0\leq u_{k,1},\dots,u_{k,m_{k}}$ $\in$ $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}$ and $0\leq y_{k,1},\dots,y_{k,m_{k}}\in F_{k}$ satisfy
$|w_{k}|\leq\sum_{j=1}^{m_{k}}z_{k,j}\otimes y_{k,j}=\sum_{j=1}^{m_{k}}u_{k,j}\otimes v_{k,j},\quad\lim_{k\to\infty}\max_{1\leq j\leq m_{k}}\|v_{k,j}\|=\lim_{k\to\infty}\frac{1}{k}=0,$

### v2

@@ 3 Regular nuclear polynomials
Since $\widehat{\otimes}_{n,s,|\pi|}E\otimes\,F$ is dense in $\widehat{\otimes}_{n,s,|\pi|}E\,\widehat{\otimes}_{|\pi|}\,F$ there exists $w_{0}=\displaystyle\sum_{j=1}^{m}u_{j}\otimes y_{j}\in\widehat{\otimes}_{n,s,|\pi|}E\otimes\,F$ such that $\|w-w_{0}\|_{|\pi|}\leq\dfrac{\varepsilon}{4}$. Without loss of generality, we assume that $u_{j}\neq 0$ and $y_{j}\neq 0$ for every $j=1,\dots,n$. Now, from the density of ${\otimes}_{n,s}E$ in $\widehat{\otimes}_{n,s,|\pi|}E$, for each $j=1,\dots,m$, there exists $u_{j,0}:=\displaystyle\sum_{i=1}^{m_{j}}\lambda_{i,j}\theta_{n}(x_{i}^{j})$ such that $\|u_{j}-u_{j,0}\|_{s,|\pi|}\leq\dfrac{\varepsilon}{4m\max\{\|y_{1}\|,\dots,\|y_{m}\|\}}$. Thus $u:=\sum_{j=1}^{m}\sum_{i=1}^{m_{j}}\lambda_{i,j}\theta_{n}(x_{i}^{j})\otimes y_{j}\in{\otimes}_{n,s}E\otimes F$ satisfies
$\|w-u\|_{|\pi|}$ $\leq\|w-w_{0}\|_{|\pi|}+\|w_{0}-u\|_{|\pi|}$
$\leq\frac{\varepsilon}{4}+\|\sum_{j=1}^{m}(u_{j}\otimes y_{j}-\sum_{i=1}^{m_{j}}\lambda_{i,j}\theta_{n}(x_{i}^{j})\otimes y_{j})\|_{|\pi|}$
$=\frac{\varepsilon}{4}+\|\sum_{j=1}^{m}(u_{j}-\sum_{i=1}^{m_{j}}\ \lambda_{i,j}\theta_{n}(x_{i}^{j}))\otimes y_{j}\|_{|\pi|}$
$\leq\frac{\varepsilon}{4}+\sum_{j=1}^{m}\|u_{j}-\sum_{i=1}^{m_{j}}\lambda_{i,j}\theta_{n}(x_{i}^{j})\|_{s,|\pi|}\|y_{j}\|\leq\frac{\varepsilon}{2}.$
Also, by the definition of the positive projective tensor norm, there exist $z_{1},\dots,z_{p}\in(\widehat{\otimes}_{n,s,|\pi|}E)^{+}$ and $v_{1},\dots,v_{p}\in F^{+}$ such that
$|u|\leq\sum_{i=1}^{p}z_{i}\otimes v_{i}\,\text{ and }\,\sum_{i=1}^{p}\|z_{i}\|_{s,|\pi|}\|v_{i}\|\leq\|u\|_{|\pi|}+\frac{\varepsilon}{4}.$
Without loss of generality, we assume that $v_{i}\neq 0$ for every $i=1,\dots,p$. For each $i=1,\dots,p$, we apply the definition of the positive projective symmetric tensor norm to obtain $z_{i}^{1},\dots,z_{i}^{p_{i}}\in E^{+}$ such that
$z_{i}=|z_{i}|\leq\sum_{k=1}^{p_{i}}\theta_{n}(z_{i}^{k})\,\text{ and }\,\sum_{k=1}^{p_{i}}\|z_{i}^{k}\|^{n}\leq\|z_{i}\|_{s,|\pi|}+\frac{\varepsilon}{4p\max\{\|v_{1}\|,\dots,\|v_{p}\|\}}.$
Therefore,
$|u|\leq\sum_{i=1}^{p}z_{i}\otimes v_{i}\leq\sum_{i=1}^{p}(\sum_{k=1}^{p_{i}}\theta_{n}(z_{i}^{k}))\otimes v_{i}=\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\theta_{n}(z_{i}^{k})\otimes v_{i}$
and
$\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\|z_{i}^{k}\|^{n}\|v_{i}\|$ $=\sum_{i=1}^{p}\left(\sum_{k=1}^{p_{i}}\|z_{i}^{k}\|^{n}\right)\|v_{i}\|\leq\sum_{i=1}^{p}\left(\|z_{i}\|_{s,|\pi|}+\frac{\varepsilon}{4p\max\{\|v_{1}\|,\dots,\|v_{p}\|\}}\right)\|v_{i}\|$
$=\sum_{k=1}^{p}\|z_{i}\|_{s,|\pi|}\|v_{i}\|+\frac{\varepsilon}{4}\leq\|u\|_{|\pi|}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}=\|u\|_{|\pi|}+\frac{\varepsilon}{2}.$
Let
$E_{0}:={\rm span}\{x_{1}^{j},\dots,x_{m_{1}}^{j},\,z_{i}^{1},\dots,z_{i}^{p_{i}}\,:j=1,\dots,m,\,i=1,\dots,p\}\subset E$
and
$F_{0}:={\rm span}\{y_{1},\dots,y_{m},\,v_{1},\dots,v_{p}\}\subset F.$
Since $E$ and $F$ are Dedekind complete, by [4, Proposition 2.1], for each $\delta>0$, there are finite-dimensional vector sublattices $E_{\delta}\subset E$, $F_{\delta}\subset F$, and linear maps
$U_{\delta}:E_{0}\to E_{\delta},\,V_{\delta}:F_{0}\to F_{\delta}$
such that
$\|U_{\delta}(x)-x\|\leq\delta\|x\|\,\text{for every $x\in E_{0}$}\,\text{ and }\,{\|V_{\delta}(y)-y\|\leq\delta\|y\|}\,\text{for every $y\in F_{0}$}.$
For each $\delta>0$, set $v^{\delta}=\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\theta_{n}(U_{\delta}(z_{i}^{k}))\otimes V_{\delta}(v_{i})$. We claim that $v^{\delta}\longrightarrow v:=\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\theta_{n}(z_{i}^{k})\otimes v_{i}$ as $\delta\longrightarrow 0$. We have
$\|v-v^{\delta}\|_{|\pi|}$ $\leq\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\|\theta_{n}(U_{\delta}(z_{i}^{k}))\otimes V_{\delta}(v_{i})-\theta_{n}(z_{i}^{k})\otimes v_{i}\|_{|\pi|}$
$\leq\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}[\|\theta_{n}(U_{\delta}(z_{i}^{k}))\|_{s,|\pi|}\|V_{\delta}(v_{i})-v_{i}\|+\|\theta_{n}(U_{\delta}(z_{i}^{k}))-\theta_{n}(z_{i}^{k})\|_{s,|\pi|}\|v_{i}\|]$
$\leq\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}[\|U_{\delta}(z_{i}^{k})\|^{n}\|V_{\delta}(v_{i})-v_{i}\|+n\,\max\{\|U_{\delta}(z_{i}^{k})\|,\|z_{i}^{k}\|\}^{n-1}\|U_{\delta}(z_{i}^{k})-z_{i}^{k}\|\|v_{i}\|]$
$\leq\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}[M^{n}\delta\|v_{i}\|+nM^{n}\delta\|z_{i}^{k}\|]$
$\leq\delta\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}[M^{n+1}+nM^{n+1}],$
provided $\delta\leq 1$, where $M=\displaystyle\max_{i=1,\dots,p}\{\|U_{\delta}(z_{i}^{1})\|,\dots,\|U_{\delta}(z_{i}^{p_{i}})\|,\|z_{i}^{1}\|,\dots,\|z_{i}^{p_{i}}\|,\|v_{i}\|\}$. Letting $\delta\to 0$, we obtain that $\|v-v^{\delta}\|_{|\pi|}\longrightarrow 0$. Analogously, we have that
$u^{\delta}:=\sum_{j=1}^{m}\sum_{i=1}^{m_{j}}\lambda_{i,j}\theta_{n}(U_{\delta}(x_{i}^{j}))\otimes V_{\delta}(y_{j})\longrightarrow u.$
The continuity of the lattice operations yield that
$w^{\delta}:=v^{\delta}\wedge((-v^{\delta})\vee u^{\delta})\to v\wedge((-v)\vee u)=u\quad\text{as }\delta\to 0.$
Now, we choose $0<\delta_{0}<\min\{1,\varepsilon\}$ such that $\|u-w^{\delta_{0}}\|_{|\pi|}\leq\frac{\varepsilon}{2}$. We will prove that $E_{\delta_{0}}$, $F_{\delta_{0}}$ and $\widetilde{w}:=w^{\delta_{0}}$ satisfy this lemma hypothesis. First, we note that since $E_{\delta_{0}}$ and $F_{\delta_{0}}$ are finite-dimensional vector lattices, $\otimes_{n,s}E_{\delta_{0}}\otimes F_{\delta_{0}}$ is finite-dimensional vector lattice, and so $\tilde{w}\in\otimes_{n,s}E_{\delta_{0}}\otimes F_{\delta_{0}}$. Secondly, we have that
$\|w-\tilde{w}\|_{|\pi|}\leq\|w-u\|_{|\pi|}+\|u-\tilde{w}\|_{|\pi|}\leq\varepsilon.$
Moreover, since
$|\widetilde{w}|\leq|v^{\delta_{0}}|\leq\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}|\theta_{n}(U_{\delta_{0}}(z_{i}^{k}))|\otimes|V_{\delta_{0}}(v_{i})|=\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\theta_{n}(|U_{\delta_{0}}(z_{i}^{k})|)\otimes|V_{\delta_{0}}(v_{i})|,$
we get
$\|\widetilde{w}\|_{\otimes_{n,s}E_{\delta_{0}}\widehat{\otimes}_{|\pi|}F_{\delta_{0}}}$ $\leq\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\|\theta_{n}(U_{\delta_{0}}(z_{i}^{k}))\|_{s,|\pi|}\|V_{\delta_{0}}(v_{i})\|$
$=\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\|U_{\delta_{0}}(z_{i}^{k})\|^{n}\|V_{\delta_{0}}(v_{i})\|\leq\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}(1+\delta_{0})^{n+1}\|z_{i}^{k}\|^{n}\|v_{i}\|$
$\leq(1+\delta_{0})^{n+1}(\|u\|_{|\pi|}+\frac{\varepsilon}{2})\leq(1+\varepsilon)^{n+1}(\|\widetilde{w}\|_{|\pi|}+\varepsilon)\leq(1+\varepsilon)^{n+1}\|\widetilde{w}\|_{|\pi|}+2^{n}\varepsilon,$
and we are done. ∎
We will also need the following technical lemma.
Lemma 3.2.
Let $E$ and $F$ be two Dedekind complete Banach lattices. For every $w\in\widehat{\otimes}_{n,s,|\pi|}E\,\widehat{\otimes}_{|\pi|}F$, there exist sequences $(E_{k})_{k}$ and $(F_{k})_{k}$ of finite-dimensional vector sublattices of $E$ and $F$, respectively, and a sequence $(w_{k})_{k}$ in $\widehat{\otimes}_{n,s,|\pi|}E\,\widehat{\otimes}_{|\pi|}F$ such that $\|w\|_{|\pi|}=\sum_{k=1}^{\infty}\|w_{k}\|_{|\pi|}<\infty$, $w=\sum_{k=1}^{\infty}w_{k}$, $w_{k}\in\otimes_{n,s}E_{k}\otimes F_{k}$ and $\|w_{k}\|_{\otimes_{n,s}E_{k}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq 2^{n}\|w_{k}\|_{|\pi|}+2^{n-k}$ for every $k\in\mathbb{N}$. Moreover, $\|w_{k}\|_{\otimes_{n,s}E_{k}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq\frac{2^{n}\cdot 3}{2^{k}}$ holds for every $k\in\mathbb{N}$ whenever $\|w\|_{|\pi|}=1$.
Proof.
@@
Let $w\in\widehat{\otimes}_{n,s,|\pi|}E\,\widehat{\otimes}_{|\pi|}F$ be given. By Lemma 3.1, there exist finite dimensional vector sublattices $E_{1}\subset E$ and $F_{1}\subset F$, and a vector $w_{1}\in\otimes_{n,s}E_{1}\,\otimes F_{1}$ such that $\|w-w_{1}\|_{|\pi|}\leq 2^{-1}$ and
$\|w_{1}\|_{\otimes_{n,s}E_{1}\,\widehat{\otimes}_{|\pi|}F_{1}}\leq\left(1+\frac{1}{2}\right)^{n}\|w_{1}\|_{|\pi|}+2^{n}\dfrac{1}{2}\leq 2^{n}\|w_{1}\|_{|\pi|}+2^{n-1}.$
Using Lemma 3.1 again, there exist finite dimensional vector sublattices $E_{2}\subset E$ and $F_{2}\subset F$, and a vector $w_{2}\in\otimes_{n,s}E_{2}\,\otimes F_{2}$ such that $\|(w-w_{1})-w_{2}\|_{|\pi|}\leq 2^{-2}$ and
$\|w_{2}\|_{\otimes_{n,s}E_{2}\,\widehat{\otimes}_{|\pi|}F_{2}}\leq\left(1+\frac{1}{2^{2}}\right)^{n}\|w_{2}\|_{|\pi|}+2^{n}\frac{1}{2^{2}}\leq 2^{n}\|w_{2}\|_{|\pi|}+2^{n-2}.$
Proceeding inductively, we may construct sequences $(E_{k})_{k}$ and $(F_{k})_{k}$ of finite dimensional vectors sublattices of $E$ and $F$, respectively, and a sequence $(w_{k})_{k}\subset\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ such that $w_{k}\in\otimes_{n,s}E_{k}\,\otimes F_{k}$, $\|(w-w_{1}-\cdots-w_{k-1})-w_{k}\|_{|\pi|}\leq 2^{-k}$ and
$\|w_{k}\|_{\otimes_{n,s}E_{k}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq 2^{n}\|w_{k}\|_{|\pi|}+2^{n-k}\,\text{ for every $k>1$.}$
Thus,
$\left|\|w\|_{|\pi|}-\sum_{i=1}^{k}\|w_{i}\|_{|\pi|}\right|\leq\left\|w-\sum_{i=1}^{k}w_{i}\right\|_{|\pi|}\leq 2^{-k}$
@@
holds for every $k\in\mathbb{N}$. By letting $k\to\infty$, we get that $\|w\|_{|\pi|}=\sum_{i=1}^{\infty}\|w_{i}\|_{|\pi|}<\infty$ and $w=\displaystyle\sum_{k=1}^{\infty}w_{k}$. We also note that $\|w_{k+1}\|_{|\pi|}\leq\|w\|_{|\pi|}-\sum_{i=1}^{k}\|w_{i}\|_{|\pi|}\leq 2^{-k}$ holds for every $k\in\mathbb{N}$, and so
$\|w_{k}\|_{\otimes_{n,s}E_{k}\,\widehat{\otimes}_{|\pi|}F_{k}}$ $\leq 2^{n}\|w_{k}\|_{|\pi|}+2^{n-k}\leq 2^{n}\,2^{-(k-1)}+2^{n-k}$
$=2^{n}\,2^{-k+1}+2^{n-k}=2^{n-k}(2+1)=\frac{2^{n}\cdot 3}{2^{k}}$
@@
$\|w_{1}\|_{\otimes_{n,s}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{1}}\leq 2^{n}\|w_{1}\|_{|\pi|}+2^{n-1}\leq 2^{n}\left(\|w\|_{|\pi|}+\frac{1}{2}\right)\leq\displaystyle\frac{2^{n}\cdot 3}{2^{1}},$
@@
Let $E$ and $F$ be Banach lattices with $F$ Dedekind complete. If $E^{*}$ and $F$ have the LAP, then the map $\overline{\Phi}_{n}\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;F)$ is injective, and hence an isometry.
@@
Suppose for the sake of contradiction that $\overline{\Phi}_{n}$ is not injective. Thus, there exists $w\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ with $\|w\|_{|\pi|}=1$ such that $\overline{\Phi}_{n}(w)=0$. Since $E^{*}$ and $F$ are Dedekind complete, we can apply Lemma 3.2 to obtain sequences $(G_{k})_{k}$ and $(F_{k})_{k}$ of finite-dimensional vector sublattices of $E^{*}$ and $F$, respectively, and the corresponding sequence $(w_{k})_{k}\subset\widehat{\otimes}_{n,s,|\pi|}E^{*}\widehat{\otimes}_{|\pi|}F$. Since, for every $k\in\mathbb{N}$, $w_{k}\in\otimes_{n,s}G_{k}\otimes F_{k}$ and $\|w_{k}\|_{\otimes_{n,s}G_{k}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq 2^{n}\|w_{k}\|_{|\pi|}+2^{n-k}\leq\displaystyle\frac{2^{n}\cdot 3}{2^{k}}$, there are a positive elements $z_{k,1},\dots,z_{k,m_{k}}\in\otimes_{n,s}G_{k}$ and positive vectors $y_{k,1},\dots,y_{k,m_{k}}\in F_{k}$ such that
$|w_{k}|\leq\displaystyle\sum_{j=1}^{m_{k}}z_{k,j}\otimes y_{k,j}\quad\text{ and}\quad\displaystyle\sum_{j=1}^{m_{k}}\|z_{k,j}\|_{s,|\pi|}\|y_{k,j}\|\leq\frac{2^{n}\cdot 3}{2^{k}}.$
Claim 1: For each $k\in\mathbb{N}$, there are $0\leq u_{k,1},\dots,u_{k,m_{k}}$ $\in$ $\otimes_{n,s}G_{k}$ and $0\leq v_{k,1},\dots,v_{k,m_{k}}\in F_{k}$ such that
$|w_{k}|\leq\sum_{j=1}^{m_{k}}u_{k,j}\otimes v_{k,j},\,\max_{1\leq j\leq m_{k}}\|v_{k,j}\|=\frac{1}{k},\,\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\|u_{k,j}\|_{s,|\pi|}<\infty.$ (1)
Indeed, first note that we can assume that $y_{k,j}\not=0$ as if it is equal to 0, then $u_{k,j}\otimes z_{k,j}=0$. For each $k\in\mathbb{N}$ and each $j=1,\dots,m_{k}$, we let
$u_{k,j}=k\|y_{k,j}\|z_{k,j}\quad\text{and}\quad v_{k,j}=\frac{y_{k,j}}{k\|y_{k,j}\|}.$
Thus, $0\leq u_{k,1},\dots,u_{k,m_{k}}$ $\in$ $\otimes_{n,s}G_{k}$ and $0\leq y_{k,1},\dots,y_{k,m_{k}}\in F_{k}$ satisfy
$|w_{k}|\leq\sum_{j=1}^{m_{k}}z_{k,j}\otimes y_{k,j}=\sum_{j=1}^{m_{k}}u_{k,j}\otimes v_{k,j},\quad\max_{1\leq j\leq m_{k}}\|v_{k,j}\|=\frac{1}{k},$